# Image inpainting for corrupted images by using the semi-super resolution GAN

### v1

@@ Abstract
Index Terms:
Deep learning, GAN, image inpainting
@@ I Introduction
Deep learning has revolutionized numerous fields, including classification [1], noise reduction [2], semantic segmentation [3], and even channel estimation in wireless communication [4]. One of the pivotal aspects of photo editing resides in the technique of image inpainting, which has undergone significant advancement through the application of deep learning methodologies. Within this domain, a central challenge is the precise reconstruction of missing image segments or pixels or the removal of undesirable elements from photographs while simultaneously reconstructing the background of the edited region. These tools serve a vital purpose in the restoration of aged photographs, striving to closely mirror the original image. Consequently, the utilisation of generative models has proven to be instrumental in achieving the highest level of fidelity in photo restoration. It should be noted that augmenting the extent of damage is an unavoidable consequence of expanding the scope of restoration.
@@ II-B Network architecture
[…]
@@
$\min_{G}\max_{D}V(D,G)=\mathbb{E}_{\textbf{H}_{\textit{SR}}\sim p_{\textit{train}}(\textbf{H}_{\textit{SR}})}[\log D(\textbf{H}_{\textit{SR}})]+\dots{\\ }\mathbb{E}_{\textbf{H}_{\textit{LR}}\sim p_{G}(\textbf{H}_{\textit{LR}})}[\log(1-D(G(\textbf{H}_{\textit{LR}})))]$ (1)
@@ II-C Loss Function
$\textit{Loss}_{D}=\mathcal{L}_{F}+\mathcal{L}_{R}$ (2)
@@
$\mathcal{L}_{R}=\textit{MSE}(\textbf{D}_{R},\textbf{1}-0.1\boldsymbol{\alpha})$ (3)
@@
$\mathcal{L}_{F}=\textit{MSE}(\textbf{D}_{F},\textbf{S}_{D})$ (4)
@@
$\textit{Loss}_{G}=\textit{MSE}(\hat{\textbf{H}}_{\textit{SR}},\textbf{H}_{\textit{SR}})+10^{-3}(\textit{MSE}(\textbf{D}_{F},\textbf{1}))$ (5)
@@ III Results
$\textit{NMSE}=\mathbb{E}_{\textbf{H}_{\textit{SR}}}\biggl{\{}\frac{||\textbf{H}_{\textit{SR}}-\hat{\textbf{H}}_{\textit{SR}}||^{2}_{F}}{||\textbf{H}_{\textit{SR}}||^{2}_{F}}\biggr{\}}$ (6)

### v2

@@ I Introduction
Deep learning has revolutionized numerous fields, including classification [17], noise reduction [11], object detection [16], and even channel estimation in wireless communication [9, 10]. One of the pivotal aspects of photo editing resides in the technique of image inpainting, which has undergone significant advancement through the application of deep learning methodologies. Within this domain, a central challenge is the precise reconstruction of missing image segments or pixels or the removal of undesirable elements from photographs while simultaneously reconstructing the background of the edited region. These tools serve a vital purpose in the restoration of aged photographs, striving to closely mirror the original image. Consequently, the utilisation of generative models has proven to be instrumental in achieving the highest level of fidelity in photo restoration. It should be noted that augmenting the extent of damage is an unavoidable consequence of expanding the scope of restoration.
@@ II-B Network architecture
The generative adversarial networks consist of two main parts that were named Generator $(\textit{G})$ and Discriminator $(\textit{D})$. The generator gives the input image a size of $(C\times H\times W)$ and the output size is the same as the input dimension.
@@
$\min_{G}\max_{D}V(D,G)=\mathbb{E}_{\textbf{H}_{\textit{SR}}\sim p_{\textit{train}}(\textbf{H}_{\textit{SR}})}[\log D(\textbf{H}_{\textit{SR}})]+\dots{\\ }\mathbb{E}_{\textbf{H}_{\textit{LR}}\sim p_{G}(\textbf{H}_{\textit{LR}})}[\log(1-D(G(\textbf{H}_{\textit{LR}})))].$ (1)
@@ II-C Loss Function
$\textit{Loss}_{D}=\mathcal{L}_{F}+\mathcal{L}_{R}.$ (2)
@@
[…]
@@
$\mathcal{L}_{F}=\textit{MSE}(\textbf{D}_{F},\textbf{S}_{D}).$ (4)
@@
[…]
@@ III Results
[…]
In addition to NMSE, we also report the Peak Signal-to-Noise Ratio (PSNR), a widely used evaluation metric in image inpainting and image restoration tasks. PSNR quantifies the reconstruction fidelity by comparing the mean squared error (MSE) between the reconstructed and ground-truth images to the maximum possible pixel intensity, and is defined as:
[…]
Since NMSE is derived from the same squared reconstruction error, PSNR, as defined in (7), can be directly related to NMSE through the mean squared error (MSE) term. While NMSE quantifies the normalized energy of the reconstruction error, PSNR expresses the same error on a logarithmic scale relative to the signal dynamic range. Consequently, the inclusion of PSNR provides a complementary evaluation perspective and facilitates direct comparison with existing image inpainting and super-resolution methods reported in the literature.
Table II presents a quantitative comparison between NMSE and PSNR values obtained at the final training epoch for all evaluated datasets. The reported results correspond to a pixel corruption level of 50%, illustrating the inverse relationship between the two metrics, where lower NMSE values consistently result in higher PSNR.
TABLE II: PSNR and NMSE
Dataset PSNR NMSE
Oxfordiiitpet 20.36 0.0092
Caltec101 21.14 0.0077
Flower102 17.99 0.0159